OpenStudy (anonymous):

Water freezes at 0 Celsius and 32 Fahrenheit. Water boils at 100 Celsius and 212 Fahrenheit. a. Write degrees Celsius as a linear function of degrees Fahrenheit . b. What is the slope of your linear equation? What does it mean? c. What is the y-intercept and what does it mean? Help and please explain. Will give medal!

OpenStudy (anonymous):

B

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

I need help with the different parts of the question.

OpenStudy (anonymous):

a linear function when you see those words think "line" as in y= mx+b notice that y depends on x. (to get y's value, you sub in a value for x and do the arithmetic) so part (a) Celsius as a linear function of degrees Fahrenheit means treat Celsius as y ( it depends on Fahrenheit) and Fahrenheit as x You have two (x,y) points: (32,0) and (212,100) can you find the equation of a line through two points? at the end, relabel x as F and y as C

OpenStudy (anonymous):

so (b) the slop = 5/9. It means that 1 degree change in Fahrenheit is only 5/9 of a degree change in Celsius. (Fahrenheit degrees are smaller)

OpenStudy (anonymous):

just tell me if it is confusing

OpenStudy (anonymous):

ok so I got the equation C=9/5+32. Is that right?

OpenStudy (anonymous):

y= 5/9 x + b

OpenStudy (anonymous):

Are you able to switch around the two points given because I got the slope of 9/5

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

because I used the slope formula with the ordered pairs of (0,32) and(100,212)

OpenStudy (anonymous):

look like your right

OpenStudy (anonymous):

So the slope would be in fact 9/5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Hold on.

OpenStudy (anonymous):

what

OpenStudy (anonymous):

so what would be the eqation?

OpenStudy (anonymous):

x=32 y=0 plug it in

OpenStudy (anonymous):

C=9/5+32

OpenStudy (anonymous):

dont forget the 0

OpenStudy (anonymous):

you dont need the zero because when you put it in the point slope formula it 32 will be added to it

OpenStudy (anonymous):

you have to find b

OpenStudy (anonymous):

it wold be zero

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

0=5/9*32+b

OpenStudy (anonymous):

there we go

OpenStudy (anonymous):

Thanks so much for your help.

OpenStudy (anonymous):

your welcome